4. What is the coefficient of performance () of a carnot refrigerator workingbetween 30°C and 0°C?

Grade:11

1 Answers

Arun
6 years ago
Dear Vaibhav
 
Coefficient of performance (COP) = T2/(T1 - T2)
 = 273/(303- 273)
 = 273/30
= 9.1
 
Regards
Arun (askIITians forum expert)